A object was fired at 17meters per second at a 40 degree and traveled 29 meters how long was it in the air


1
Expert's answer
2021-10-27T08:48:17-0400

Let's first find the time that the object takes to reach the maximum height:


"v_y = v_{0y}-gt,""0=v_0sin\theta-gt,""t=\dfrac{v_0sin\theta}{g}."


Then, we can find the flight time of the object (how long was it in the air) as follows:


"t_{flight}=2t=\dfrac{2v_0sin\theta}{g},""t_{flight}=\dfrac{2\times17\ \dfrac{m}{s}\times sin40^{\circ}}{9.8\ \dfrac{m}{s^2}}=2.23\ s."